# When Do Neural Networks Learn World Models?

### v4

@@ 2 Formulation of Learning World Models
${\mathcal{H}}(h)\vcentcolon=\{f:{\mathcal{X}}\to{\mathbb{R}}\mid$ $f({\bm{x}})=h({\bm{x}}),$ (1)
@@ 4.1 Low-Degree Bias Drives Representation Learning
Intuitively, this result is a consequence of learning “redundant” representations: for any $h$, it suffices to learn every parity function $\chi_{S}$ with $\hat{h}(S)\neq 0$ in its Fourier-Walsh transform. Even if there is a good universal data representation, explicitly learning it is often not the best choice since it may involve irrelevant parity functions with $\hat{h}(S)=0$, resulting in a larger realization degree.
@@ 4.3 Benefits of Learning World Models
Let the latent variables during training be uniformly sampled from the Hamming ball $B_{r}\vcentcolon=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$ with $r<d$, and let the latent variables in test be uniformly sampled from ${\mathcal{Z}}$. Let $h:\{\pm 1\}^{m}\to\{\pm 1\}$ be a downstream task such that $h\circ\psi$ is a parity function with $\mathrm{deg}(h)=q>k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose i}\rceil$. Then, if $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and $\mathrm{deg}(h\circ\psi)>k-q+r$, the following holds: (i) the test mean square error (MSE) of any $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$ is larger than $1$; (ii) let $\Phi^{*}$ be a representation that learns the world model up to negations and permutations as in Theorem 4.8 and let $g^{*}$ be a function such that $g^{*}\circ\Phi^{*}\in{\mathcal{H}}(h)$, then the test MSE of $g^{*}\circ\Phi^{*}$ is 00.
@@ 4.4 Impact of Model Architecture
Notably, for a given input dimension $n$, the Fourier-Walsh transform uses parity functions $\chi_{S}$ as a basis of the $2^{n}$-dimensional vector space ${\mathcal{F}}^{n}=\{f:\{\pm 1\}\to{\mathbb{R}}\}$ (cf. Definition 3.1). To obtain a different basis, we define a basis transform $U$, i.e., an invertible linear transform on ${\mathcal{F}}^{n}$ such that $\{U(\chi_{S})\mid S\subseteq[n]\}$ is a basis of ${\mathcal{F}}^{n}$. The degree of a function $f:\{\pm 1\}^{n}\to{\mathbb{R}}$ under the new basis is then given by
@@
where $U^{-1}$ reflects the cost of using the new basis to represent the original one. As a result, the low-degree bias under $\mathrm{deg}_{U}(\cdot)$ may deviate from that under $\mathrm{deg}(\cdot)$. To capture the effect of this, we introduce the notion of basis compatibility.
@@
We say a basis transform $U$ is compatible if $\mathrm{deg}(U(\chi_{S}))=\mathrm{deg}(\chi_{S}),\,\forall S\subseteq[n],\forall n$.
@@
Consider the same setting as in Theorem 4.8 with $n\to\infty$. Let $U$ be a basis transform and let the degrees of $\Phi$ and $g$ be measured under the new basis $\{U(\chi_{S})\}$. Then, (i) if $U$ is compatible, $\Phi^{*}$ learns the world model up to negations and permutations; (ii) there exists incompatible $U$ such that $\Phi^{*}=T\circ\psi^{-1}$, where $T$ is an invertible transform on ${\mathcal{Z}}$ satisfying $\max_{i\in[d]}\mathrm{deg}(T^{-1}_{i})\geq k$.
@@ Appendix B Preliminaries on Boolean Function Analysis
There are different ways to represent the input bits in the above definiton. A natural way is to use 00 and $1$ as elements of the field ${\mathbb{F}}_{2}$. In this way, a (single-output) Boolean function $f$ with $n$ input coordinates (bits) is represented by $f:\{0,1\}^{n}\to\{0,1\}$. It is also convenient to use $-1$ and $1$, thought as real numbers, and define $f$ as a function from $\{-1,1\}^{n}$ to $\{-1,1\}$. The latter representation can be easily transformed from the former one by the mapping $b\mapsto(-1)^{b}$ over $\{0,1\}$. In our analysis, we will mostly use the latter representation as it is more compatible with the Fourier-Walsh transform of Boolean functions.
@@
It is easy to verify that ${\bm{1}}_{\bm{a}}({\bm{x}})$ takes value $1$ when ${\bm{x}}={\bm{a}}$ and 00 otherwise. We then have
@@
[…]
@@ Appendix C Discussion on Complexity Measures
For example, the linear function $f(x)=ax+b,x\in{\mathbb{R}}$ does not have a standard Fourier transform since it is not absolutely integrable. Even if we constrain it to some finite interval $[-L,L]$ and periodically extend it across the real axis, the resulting Fourier transform would involve frequencies up to infinity, suggesting infinite complexity if we use the highest frequency as the complexity measure. This is in contrast to the empirically observed simplicity bias of gradient descent, where linear models are often learned first (see e.g., Kalimeris et al. (2019)). The measure of polynomial order has a similar problem since not all continuous functions can be linearly represented by a set of polynomials. In fact, the degree of Boolean functions could be viewed as a “discrete” version of polynomial order, but it is more principled since the Boolean function space indeed has a polynomial basis given by the Fourier-Walsh transform. Finally, we found other compression-based approximations of Kolmogorov complexity hard to analyze theoretically since they rely on specific compression algorithms.
@@ Appendix D Numerical Experiments
This section presents numerical experiments that corroborate our theoretical results. In all experiments, we train MLPs using stochastic gradient descent (SGD) and binary data from {±$\{\pm 1\}^{m}$, generated by the data generation model in Table 1 with $d=m=10$. A summarization of results is as follows:
@@
Figure 3: Low-degree bias of neural networks: Training latents are uniformly sampled from the Hamming ball $B_{r}=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$; the out-of-distribution (OOD) test latents are uniformly sampled from {±$\{\pm 1\}^{d}\,(d=10)$. For small $r$, low-degree realizations of tasks (target functions) exist in the training distribution. A model that learns these low-degree realizations would have low in-distribution (ID) error yet high OOD error due to the loss of high-degree components in target functions. We observe that while SGD-trained MLPs learn ground-truth target functions when $r=10$ (denoted by “Full” in the legends), they indeed fit low-degree realizations for training data with $r=2$ and $r=3$, resulting in high OOD error. We separately train a three-layer MLP for each task and for each $r$.
(a) $2$-degree tasks
(b) $3$-degree tasks
(c) $4$-degree tasks
@@ Appendix E Proofs
  • In Section E.4, we provide the proof of Theorem 4.4.
@@ E.1 Technical Lemmas
holds for some $b_{1},\ldots,b_{d+1}$, where $b_{S^{\prime}}\in\{-1,1\}$ for every $S^{\prime}\subset S$. This means that we can thus replace every degree-$d+1$ monomial $\chi_{S}({\bm{x}})=\prod_{j\in S}x_{j}$ by a degree-$d$ polynomial $-\sum_{S^{\prime}\in 2^{S},|S^{\prime}|\leq d}b_{S^{\prime}}\prod_{k\in S^{\prime}}x_{k}$. By iteratively using this replacement in the Fourier-Walsh transform of $h$ (Definition 3.1), one can eventually obtain a polynomial with degree $d$ or less without changing the value of the function on every ${\bm{x}}\in{\mathcal{X}}$. This completes the proof. ∎
@@
Let $f:\{-1,1\}^{n}\to\{-1,1\}^{n}$ be a bijective function and let $11\leq k\leq n-$ 1 be an integer. Then, we have $\mathrm{deg}(f_{i})=1,\forall i\in[n]$ if and only if for every $S\subset[n]$ with $|S|=k$, there exists $T\subset[n]$ with $|T|=k$ such that:
@@
Note that $\mathrm{deg}(f_{i})=1,\forall i\in[n]$ together with the fact that $f$ is bijective indicates the existence of a permutation $i_{1},\ldots,i_{n}$ of $1,\ldots,n$ such that $f_{j}({\bm{x}})=x_{i_{j}}$ or $f_{j}({\bm{x}})=-x_{i_{j}}$ for every $j\in[n]$, which trivially gives the result. In the following we prove the other direction. Note that the case of $k=1$ is trivial. Hence, to prove that $\mathrm{deg}(f_{i})=1$, it suffices to prove (*): for every $i\in[n]$, there exists $j\in[n]$ such that $\mathrm{Inf}_{i}(f_{j})>0$ and $\mathrm{Inf}_{i}(f_{m})=0,\forall m\neq j,m\in[n]$, for every $12\leq k\leq n-$ 1.
@@
We first prove that for every $S$, $T$ is unique, by contradiction. Suppose that for some $S$, $T$ is not unique, i.e., there exists $T^{\prime}\neq T\subset[n]$ such that $T^{\prime}$ satisfies the condition. Then, by Definition E.4, changing the values of the coordinates $x_{i},i\in S$ can change only the values of $f_{j}({\bm{x}}),j\in T\cap T^{\prime}$ but not the values of other $f_{j}({\bm{x}}),j\in[n]\setminus(T\cup T^{\prime})$. This results in |$|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{n-|S|}\cdot 2^{|T\cap T^{\prime}|}=2^{n-k+|T\cap T^{\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and $T$ is unique. This allows us to define a mapping $\phi:S\mapsto T$.
@@
Let ${\mathcal{S}}_{k}=\{S\subset[n]\mid|S|=k\}$ be the set of all subsets of $[n]$ with cardinality $k$. We then prove that $\phi:S\mapsto T$ is bijective on ${\mathcal{S}}_{k}$. To this end, it suffices to show that every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$. This can be similarly proved by contradiction as in proving the uniqueness of $T$: if it is false, then there exists a subset $S^{\prime\prime}=S\cup S^{\prime}$ and $T$ such that $\mathrm{Inf}_{i}(f_{j})=0,\,\forall i\in S^{\prime\prime},\,j\in[n]\setminus T$. Then, by Definition E.4, changing the values of the coordinates $x_{i},i\in S^{\prime\prime}$ can change only the values of $f_{j}({\bm{x}}),j\in T$ but not the values of other $f_{j}({\bm{x}}),j\notin T$. This results in |$|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{|T|}\cdot 2^{n-|S^{\prime\prime}|}=2^{n+k-|S^{\prime\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$.
@@
Given that $\phi:S\mapsto T$ is bijective, we know that for every subset ${\mathcal{S}}\subseteq{\mathcal{S}}_{k}$, there exists a unique subset ${\mathcal{T}}=\{T=\phi(S)\mid S\in{\mathcal{S}}\}\subseteq{\mathcal{S}}_{k}$ such that $|{\mathcal{S}}|=|{\mathcal{T}}|$. We then move on to prove the proposition (*) by contradiction: suppose it is false, i.e., for some $i\in[n]$, there exists $U\subset[n]$ with $|U|\geq 2$ such that $\mathrm{Inf}_{i}(f_{j})>0$ for every $j\in U$ and $\mathrm{Inf}_{i}(f_{j})=0$ for every $j\in[n]\setminus U$. If $|U|>k$, then for any $S$ such that $i\in S$, there does not exist a feasible $T$, which is a contradiction. If $2\leq|U|\leq k$, consider ${\mathcal{S}}=\{S\subset{\mathcal{S}}_{k}\mid i\in S\}$ with $|{\mathcal{S}}|=C_{n-1}^{k-1}$. By the definition of $\phi$, we know that ${\mathcal{T}}\subseteq\{T\subset{\mathcal{S}}_{k}\mid U\subseteq T\}$, which gives $|{\mathcal{T}}|\leq C_{n-|U|}^{k-|U|}<C_{n-1}^{k-1}=|{\mathcal{S}}|$. Thus, the assumption is false and proposition (*) is true. This completes the proof. ∎
@@
For every ${\mathcal{G}}\subseteq{\mathcal{F}}^{d}$, let ${\mathcal{G}}\circ T=\{h^{\prime}\circ T\mid h^{\prime}\in{\mathcal{G}}\}$. By Lemma E.7, we know that the mapping $h^{\prime}\mapsto h^{\prime}\circ T$ is bijective on ${\mathcal{F}}^{d}$. We thus have $|{\mathcal{F}}^{d}_{k}\circ T|=|{\mathcal{F}}^{d}_{k}|$ for every $k$. For ${\mathcal{F}}^{d}_{k}\circ T$, there are two possibilities:
@@
$\lim_{n\to\infty}\frac{1}{n}\,\sum_{i\in[n]}\mathrm{deg}(g_{i})$ $=\mathbb{E}_{k\sim\mathrm{Discrete}(p_{1},\ldots,p_{d})}\mathbb{E}_{h\sim U({\mathcal{F}}_{k}^{d}\circ\psi^{-1})}\mathrm{deg}(h\circ f\circ T^{-1})$ (57)
@@
A limitation of Theorem 4.8 is that we requires a non-zero probability of explicitly sampling degree-$1$ tasks (i.e., $p_{1}>0$), in which latent variables are essentially observed as task outputs. If $p_{1}=0$, we can still prove equation (61) by applying Lemma E.10; in other words, we can still prove that every representation $\Phi^{*}$ that learns the world model up to negations and permutations is a minimizer of the optimization problem (10). However, such minimizers may not be unique, because Lemma E.10 only proves the equivalence between the equality and $\mathrm{deg}(T_{i})=1,\forall i\in[d]$ when $k=1$ but not $11<k\leq d-$ 1. In fact, we can construct hard examples showing that in some cases, there indeed exists other $\Phi$ that minimizes $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime}\circ T)$ for every $k\in[d]\setminus\{1\}$.
@@
Let $k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose r}\rceil$. Due to the invertibility of $\psi$, we know that |$|\{{\bm{x}}\mid p(\psi^{-1}({\bm{x}}))>0\}|=|B_{r}|=\sum_{i=0}^{r}{d\choose r}$. Therefore, for any $S=\{i_{1},\ldots,i_{k^{\prime}}\}\subseteq[d]$ such that $k^{\prime}>k$, there must exist $b_{1},\ldots,b_{k^{\prime}}\in\{-1,1\}^{k^{\prime}}$ such that
@@
for every ${\bm{x}}\in{\mathcal{X}}^{\prime}\vcentcolon=\{{\bm{x}}\mid p(\psi^{-1}({\bm{x}}))>0\}$, where $b_{S^{\prime}}\in\{-1,1\}$ for every $S^{\prime}\subset S$–if this does not hold, then we have $|{\mathcal{X}}^{\prime}|\geq 2^{k^{\prime}}>2^{k}\geq\sum_{i=0}^{r}{d\choose r}$, which contradicts $|{\mathcal{X}}^{\prime}|=\sum_{i=0}^{r}{d\choose r}$. By equation (64), we can replace every degree-$k^{\prime}$ monomial $\chi_{S}({\bm{x}})=\prod_{j\in S}x_{j}$ by a degree-$k^{\prime}-1$ polynomial $-\sum_{S^{\prime}\subset S,|S^{\prime}|\leq k^{\prime}-1}b_{S^{\prime}}\prod_{k\in S^{\prime}}x_{k}$. Iteratively using this replacement in the Fourier-Walsh transform of $h$ gives the desired result. ∎

### v5

@@ 2 Formulation of Learning World Models
${\mathcal{H}}(h)=\{f:{\mathcal{X}}\to{\mathbb{R}}\mid$ $f({\bm{x}})=h({\bm{x}}),$ (1)
@@ 4.1 Low-Degree Bias Drives Representation Learning
Intuitively, this result is a consequence of learning “redundant” representations: for every $h$, it suffices to learn every parity function $\chi_{S}$ with $\hat{h}(S)\neq 0$ in its Fourier-Walsh transform. Even if there is a good universal data representation, explicitly learning it is often not the best choice since it may involve irrelevant parity functions with $\hat{h}(S)=0$, resulting in a larger realization degree.
@@ 4.3 Benefits of Learning World Models
Let the latent variables during training be uniformly sampled from the Hamming ball $B_{r}\vcentcolon=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$ with $r<d$, and let the latent variables in test be uniformly sampled from $\mathcal{Z}$}. Let $h:\{\pm 1\}^{m}\to\{\pm 1\}$ be a downstream task such that $h\circ\psi$ is a parity function with $\mathrm{deg}(h)=q>k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose i}\rceil$. Then, if $\mathrm{deg}(h\mid\psi^{-1})\geq q-r$ and $\mathrm{deg}(h\circ\psi)>k-q+r$, the following holds: (i) the test mean square error (MSE) of any $h^{*}\in\mathcal{H}_{\mathrm{min}}(h)$ is larger than $1$; (ii) let $\Phi^{*}$ be a representation that learns the world model up to negations and permutations as in Theorem 4.8 and let $g^{*}$ be a function such that $g^{*}\circ\Phi^{*}\in{\mathcal{H}}(h)$, then the test MSE of $g^{*}\circ\Phi^{*}$ is 0.
@@ 4.4 Impact of Model Architecture
For a given input dimension $m$, the Fourier-Walsh transform uses parity functions $\chi_{S}$ as a basis of the $2^{m}$-dimensional vector space ${\mathcal{F}}^{m}=\{f:\{\pm 1\}^{m}\to{\mathbb{R}}\}$ (cf. Definition 3.1). To obtain a different basis, we define a basis transform $U$, i.e., an invertible linear transform on ${\mathcal{F}}^{m}$ such that $\{U(\chi_{S})\mid S\subseteq[m]\}$ is a basis of ${\mathcal{F}}^{m}$. The degree of a function $f:\{\pm 1\}^{m}\to{\mathbb{R}}$ under this new basis is then given by
@@
where $U^{-1}$ reflects the cost of using the new basis to represent the original one. As a result, the low-degree bias under $\mathrm{deg}_{U}(\cdot)$ may deviate from that under $\mathrm{deg}(\cdot)$. To capture the effect of such a basis transform, we introduce the notion of basis compatibility.
@@
We say a basis transform $U$ is compatible if $\mathrm{deg}(U(\chi_{S}))=\mathrm{deg}(\chi_{S}),\,\forall S\subseteq[m]$.
@@
Consider the same setting as in Theorem 4.8 with $n\to\infty$. Let $U$ be a basis transform and let the degrees of $\Phi$ and $g$ be measured under the new basis $\{U(\chi_{S})\}$. Then, (i) if $U$ is compatible, $\Phi^{*}$ learns the world model up to negations and permutations; (ii) if $U$ is not compatible, then $\Phi^{*}$ does not learn the world model up to degree-$1$ transforms and for every $k\in[d]$, there exists an incompatible $U$ such that $\Phi^{*}=T\circ\psi^{-1}$, where $T$ is an invertible transform on $\mathcal{Z}$} satisfying $\max_{i\in[d]}\mathrm{deg}(T^{-1}_{i})\geq k$.
@@ Appendix B Preliminaries on Boolean Function Analysis
There are different ways to represent the input bits in the above definiton. A natural way is to use 0 and $1$ as elements of the field ${\mathbb{F}}_{2}$. In this way, a (single-output) Boolean function $f$ with $n$ input coordinates (bits) is represented by $f:\{0,1\}^{n}\to\{0,1\}$. It is also convenient to use 1-1 and $1$, thought as real numbers, and define $f$ as a function from $\{-1,1\}^{n}$ to $\{-1,1\}$. The latter representation can be easily transformed from the former one by the mapping $b\mapsto(-1)^{b}$ over $\{0,1\}$. In our analysis, we will mostly use the latter representation as it is more compatible with the Fourier-Walsh transform of Boolean functions.
@@
It is easy to verify that ${\bm{1}}_{\bm{a}}({\bm{x}})$ takes value $1$ when ${\bm{x}}={\bm{a}}$ and 0 otherwise. We then have
@@
We define the inner product $\langle\cdot,\cdot\rangle$ of functions $f:\{-1,1\}^{n}\to{\mathbb{R}}$ and $g:\{-1,1\}^{n}\to{\mathbb{R}}$ by
@@ Appendix C Discussion on Complexity Measures
For example, the linear function $f(x)=ax+b,x\in{\mathbb{R}}$ does not have a standard Fourier transform since it is not absolutely integrable. Even if we constrain it to some finite interval [L,L][-L,L] and periodically extend it across the real axis, the resulting Fourier transform would involve frequencies up to infinity, suggesting infinite complexity if we use the highest frequency as the complexity measure. This is in contrast to the empirically observed simplicity bias of gradient descent, where linear models are often learned first (see e.g., Kalimeris et al. (2019)). The measure of polynomial order has a similar problem since not all continuous functions can be linearly represented by a set of polynomials. In fact, the degree of Boolean functions could be viewed as a “discrete” version of polynomial order, but it is more principled since the Boolean function space indeed has a polynomial basis given by the Fourier-Walsh transform. Finally, we found other compression-based approximations of Kolmogorov complexity hard to analyze theoretically since they rely on specific compression algorithms.
@@ Appendix D Numerical Experiments
This section presents numerical experiments that corroborate our theoretical results. In all experiments, we train MLPs using stochastic gradient descent (SGD) and binary data from $\{\pm 1\}^{m}$, generated by the data generation model in Table 1 with $d=m=10$. A summarization of results is as follows:
@@
Figure 3: Low-degree bias of neural networks: Training latents are uniformly sampled from the Hamming ball $B_{r}=\{{\bm{z}}\in\{\pm 1\}^{d}\mid\#_{-1}({\bm{z}})\leq r\}$; the out-of-distribution (OOD) test latents are uniformly sampled from $\{\pm 1\}^{d}\,(d=10)$. For small $r$, low-degree realizations of tasks (target functions) exist in the training distribution. A model that learns these low-degree realizations would have low in-distribution (ID) error yet high OOD error due to the loss of high-degree components in target functions. We observe that while SGD-trained MLPs learn ground-truth target functions when $r=10$ (denoted by “Full” in the legends), they indeed fit low-degree realizations for training data with $r=2$ and $r=3$, resulting in high OOD error. We separately train a three-layer MLP for each task and for each $r$.
(a) 22-degree tasks
(b) 33-degree tasks
(c) 44-degree tasks
@@ Appendix E Proofs
  • In Section E.4, we provide the proof of Theorem $4.$
@@ E.1 Technical Lemmas
holds for some $b_{1},\ldots,b_{d+1}$, where $b_{S^{\prime}}\in\{-1,1\}$ for every $S^{\prime}\subset S$. This means that we can thus replace every degree-$d+1$ monomial $\chi_{S}({\bm{x}})=\prod_{j\in S}x_{j}$ by a degree-$d$ polynomial $\sum_{S^{\prime}\in 2^{S},|S^{\prime}|\leq d}b_{S^{\prime}}\prod_{k\in S^{\prime}}x_{k}$. By iteratively using this replacement in the Fourier-Walsh transform of $h$ (Definition 3.1), one can eventually obtain a polynomial with degree $d$ or less without changing the value of the function on every ${\bm{x}}\in{\mathcal{X}}$. This completes the proof. ∎
@@
Let $f:\{-1,1\}^{n}\to\{-1,1\}^{n}$ be a bijective function and let $1\leq k\leq n-1$ be an integer. Then, we have $\mathrm{deg}(f_{i})=1,\forall i\in[n]$ if and only if for every $S\subset[n]$ with $|S|=k$, there exists $T\subset[n]$ with $|T|=k$ such that:
@@
Note that $\mathrm{deg}(f_{i})=1,\forall i\in[n]$ together with the fact that $f$ is bijective indicates the existence of a permutation $i_{1},\ldots,i_{n}$ of $1,\ldots,n$ such that $f_{j}({\bm{x}})=x_{i_{j}}$ or $f_{j}({\bm{x}})=-x_{i_{j}}$ for every $j\in[n]$, which trivially gives the result. In the following we prove the other direction. Note that the case of $k=1$ is trivial. Hence, to prove that $\mathrm{deg}(f_{i})=1$, it suffices to prove (*): for every $i\in[n]$, there exists $j\in[n]$ such that $\mathrm{Inf}_{i}(f_{j})>0$ and $\mathrm{Inf}_{i}(f_{m})=0,\forall m\neq j,m\in[n]$, for every $2\leq k\leq n-1$.
@@
We first prove that for every $S$, $T$ is unique, by contradiction. Suppose that for some $S$, $T$ is not unique, i.e., there exists $T^{\prime}\neq T\subset[n]$ such that $T^{\prime}$ satisfies the condition. Then, by Definition E.4, changing the values of the coordinates $x_{i},i\in S$ can change only the values of $f_{j}({\bm{x}}),j\in T\cap T^{\prime}$ but not the values of other $f_{j}({\bm{x}}),j\in[n]\setminus(T\cup T^{\prime})$. This results in $|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{n-|S|}\cdot 2^{|T\cap T^{\prime}|}=2^{n-k+|T\cap T^{\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and $T$ is unique. This allows us to define a mapping $\phi:S\mapsto T$.
@@
Let ${\mathcal{S}}_{k}=\{S\subset[n]\mid|S|=k\}$ be the set of all subsets of $[n]$ with cardinality $k$. We then prove that $\phi:S\mapsto T$ is bijective on ${\mathcal{S}}_{k}$. To this end, it suffices to show that every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$. This can be similarly proved by contradiction as in proving the uniqueness of $T$: if it is false, then there exists a subset $S^{\prime\prime}=S\cup S^{\prime}$ and $T$ such that $\mathrm{Inf}_{i}(f_{j})=0,\,\forall i\in S^{\prime\prime},\,j\in[n]\setminus T$. Then, by Definition E.4, changing the values of the coordinates $x_{i},i\in S^{\prime\prime}$ can change only the values of $f_{j}({\bm{x}}),j\in T$ but not the values of other $f_{j}({\bm{x}}),j\notin T$. This results in $|\{f({\bm{x}}),{\bm{x}}\in{\mathcal{X}}\}|\leq 2^{|T|}\cdot 2^{n-|S^{\prime\prime}|}=2^{n+k-|S^{\prime\prime}|}<2^{n}$, contradicting the bijectivity of $f$. Therefore, the assumption is false and every $S\neq S^{\prime}\in{\mathcal{S}}_{k}$ are mapped to different $T$.
@@
Given that $\phi:S\mapsto T$ is bijective, we know that for every subset ${\mathcal{S}}\subseteq{\mathcal{S}}_{k}$, there exists a unique subset ${\mathcal{T}}=\{T=\phi(S)\mid S\in{\mathcal{S}}\}\subseteq{\mathcal{S}}_{k}$ such that $|{\mathcal{S}}|=|{\mathcal{T}}|$. We then move on to prove the proposition (*) by contradiction: suppose it is false, i.e., for some $i\in[n]$, there exists $U\subset[n]$ with $|U|\geq 2$ such that $\mathrm{Inf}_{i}(f_{j})>0$ for every $j\in U$ and $\mathrm{Inf}_{i}(f_{j})=0$ for every $j\in[n]\setminus U$. If $|U|>k$, then for every $S$ such that $i\in S$, there does not exist a feasible $T$, which is a contradiction. If $2\leq|U|\leq k$, consider ${\mathcal{S}}=\{S\subset{\mathcal{S}}_{k}\mid i\in S\}$ with $|{\mathcal{S}}|=C_{n-1}^{k-1}$. By the definition of $\phi$, we know that ${\mathcal{T}}\subseteq\{T\subset{\mathcal{S}}_{k}\mid U\subseteq T\}$, which gives $|{\mathcal{T}}|\leq C_{n-|U|}^{k-|U|}<C_{n-1}^{k-1}=|{\mathcal{S}}|$. Thus, the assumption is false and proposition (*) is true. This completes the proof. ∎
@@
For every ${\mathcal{G}}\subseteq{\mathcal{F}}^{d}$, let ${\mathcal{G}}\circ T=\{h^{\prime}\circ T\mid h^{\prime}\in{\mathcal{G}}\}$. By Lemma E.7, we know that the mapping $h^{\prime}\mapsto h^{\prime}\circ T$ is bijective on ${\mathcal{F}}^{d}$. We thus have $|{\mathcal{F}}^{d}_{k}\circ T|=|{\mathcal{F}}^{d}_{k}|$ for every $k$. For ${\mathcal{F}}^{d}_{k}\circ T$, there are two possible cases:
@@
$\lim_{n\to\infty}\frac{1}{n}\,\sum_{i\in[n]}\mathrm{deg}(g_{i})$ $=\mathbb{E}_{k\sim\mathrm{Discrete}(p_{1},\ldots,p_{d})}\mathbb{E}_{h\sim U({\mathcal{F}}_{k}^{d}\circ\psi^{-1})}\mathrm{deg}(h\circ\psi\circ T^{-1})$ (57)
@@
A limitation of Theorem 4.8 is that we requires a non-zero probability of explicitly sampling degree-$1$ tasks (i.e., $p_{1}>0$), in which latent variables are essentially observed as task outputs. If $p_{1}=0$, we can still prove equation (61) by applying Lemma E.10; in other words, we can still prove that every representation $\Phi^{*}$ that learns the world model up to negations and permutations is a minimizer of the optimization problem (10). However, such minimizers may not be unique, because Lemma E.10 only proves the equivalence between the equality and $\mathrm{deg}(T_{i})=1,\forall i\in[d]$ when $k=1$ but not $1<k\leq d-1$. In fact, we can construct hard examples showing that in some cases, there indeed exists other $\Phi$ that minimizes $\sum_{h^{\prime}\in{\mathcal{F}}^{d}_{k}}\mathrm{deg}(h^{\prime}\circ T)$ for every $k\in[d]\setminus\{1\}$.
@@
Let $k=\lceil\log_{2}\sum_{i=0}^{r}{d\choose r}\rceil$. Due to the invertibility of $\psi$, we know that $|\{{\bm{x}}\mid p(\psi^{-1}({\bm{x}}))>0\}|=|B_{r}|=\sum_{i=0}^{r}{d\choose r}$. Therefore, for every $S=\{i_{1},\ldots,i_{k^{\prime}}\}\subseteq[d]$ such that $k^{\prime}>k$, there must exist $b_{1},\ldots,b_{k^{\prime}}\in\{-1,1\}^{k^{\prime}}$ such that
@@
for every ${\bm{x}}\in{\mathcal{X}}^{\prime}\vcentcolon=\{{\bm{x}}\mid p(\psi^{-1}({\bm{x}}))>0\}$, where $b_{S^{\prime}}\in\{-1,1\}$ for every $S^{\prime}\subset S$–if this does not hold, then we have $|{\mathcal{X}}^{\prime}|\geq 2^{k^{\prime}}>2^{k}\geq\sum_{i=0}^{r}{d\choose r}$, which contradicts $|{\mathcal{X}}^{\prime}|=\sum_{i=0}^{r}{d\choose r}$. By equation (64), we can replace every degree-$k^{\prime}$ monomial $\chi_{S}({\bm{x}})=\prod_{j\in S}x_{j}$ by a degree-$k^{\prime}-1$ polynomial $\sum_{S^{\prime}\subset S,|S^{\prime}|\leq k^{\prime}-1}b_{S^{\prime}}\prod_{k\in S^{\prime}}x_{k}$. Iteratively using this replacement in the Fourier-Walsh transform of $h$ gives the desired result. ∎